OpenStudy (anonymous):

Which equation is a point-slope form of the equation of this line? the points are (-2,1) and (1,7) A) −2x+y+5=0 B) y=2x−5 C) (y+1)=−12(x−2) D) (y−1)=2(x+2)

OpenStudy (muzzack):

7-1/1+2 6/3 m=2 y-1=2(x+2) y=2x+5